OpenStudy (anonymous):

What is the length of the hypotenuse of a right triangle if each of the two legs is 7 units? Square root of 14 7 Square root of 98 49

Nnesha (nnesha):

Pythagorean theorem \[\huge\rm a^2+b^2=c^2\] c is the longest side of riaght triangle which is hypotenuse plug in a ,b values solve for c

OpenStudy (anonymous):

What? That really just confused me! Lol! I'm trying to understand this, but I don't get it!

OpenStudy (anonymous):

I have nothing to plug in

Nnesha (nnesha):

to find 3rd side of *right* triangle you have to use Pythagorean theorm a^2+b^2=c^2

Nnesha (nnesha):

read the question